Graph the linear system and estimate the solution.
y=-3x+2
3x+y=5
Thank you.

Answer by mananth(16946)

y=-3x+2
3x+y=5

equation1
when x= 0 y= 2 ( 0 , 2 )
when y = 0 x= 0.67 ( 0.67 , 0 )
Equation 2
when x= 0 y= 5 ( 0 , 5 ) 5
when y = 0 x= 1.67 ( 1.67 , 0 )



They are parallel lines and there is no solution to the system